Grandma is 60 years old and her grandchild is 4 years old.
When will the grandma be 8 times older than the grandchild?
When will grandma be 5 times as old as grandchild?
When will grandma be only 3 times as old as the grandchild?

Answer by ankor@dixie-net.com(22740)

Grandma is 60 years old and her grandchild is 4 years old.
:
When will the grandma be 8 times older than the grandchild?
let a = no. of years for this to be true
a + 60 = 8(a+4)
a + 60 = 8a + 32
60 - 32 = 8a - a
28 = 7a
a = 28/7
a = 4 yrs she will be 8 times as old; (64 and 8)
:
When will grandma be 5 times as old as grandchild?
let b = no. of years for this to be true
b + 60 = 5(b+4)
b + 60 = 5b + 20
60 - 20 = 5b - b
40 = 4b
b = 40/4
b = 10 yrs she will be 5 times as old (70 and 14)
:
When will grandma be only 3 times as old as the grandchild
let c = no. of yrs for this to be true
You should be able to handle this last one by yourself